Obtain the conditions for observing maxima and minima in a Young two-slit
interference pattern. Show that these conditions change when a thin transparent
sheet of thickness t and refractive index μ is introduced in the path of one of the
superposing beams.
1
Expert's answer
2019-03-29T11:03:28-0400

1 Obtain the conditions for observing maxima and minima in a Young two-slit interference pattern.

Let S1 and S2 be two slits separated by a distance d. Consider a point P on XY plane such that CP = x. The nature of interference between two waves reaching point P depends on the path difference S2P-S1P.

We using figure:

"S1P^2=D^2+(x-\frac{d}{2})^2 (1)""S2P^2=D^2+(x+\frac{d}{2})^2 (2)"

We get using (1) and (2):


"S2P^2 - S1P^2=2xd (3)"

"(S2P - S1P)= \frac{2xd}{ S2P + S1P } (4)"

for x, d<<< D , S1P+S2P =2D with negligible error included , path difference would be


"(S2P - S1P)= \frac{xd}{D} (5)"

Phase difference between wave for constructive interference is equal to


"n \lambda"

In this case, we can write



"n \lambda=\frac{xd}{D} (6)"

where

"n=0, \pm1, \pm2, \pm3, ..."

Phase difference between wave for destructive interference is equal to


"\frac{(2n+1)\lambda }{2}"

Similarly, for destructive interference,


"\frac{(2n+1)\lambda }{2} =\frac{xd}{D} (7)"




2 Show that these conditions change when a thin transparent sheet of thickness t and refractive index μ is introduced in the path of one of the superposing beams.


Let a thin transparent sheet of thickness t and refractive index μ be introduced in the path of wave from one slit S1. It is seen from the figure that light reaching the point P from source S1 has to traverse a distance t in the sheet and a distance (S1P−t) in the air. If c and v are velocities of light in air and in transparent sheet respectively, then the time taken by light to reach from S1 to P is given by



"\frac{(S1P-t) } {c } + \frac{t} {v}=\frac{(S1P-t) } {c } + \frac{\mu t} {c} (8)"

The effective path difference at any point P on the screen



"Δ= S2P - (S1P+(\mu-1)t) (9)"


Using (5) and (9) we can write for constructive interference


"n \lambda=\frac{xd}{D} - (\mu-1)t (10)"

for destructive interference


"\frac{(2n+1)\lambda }{2}=\frac{xd}{D} - (\mu-1)t (11)"